# Federated Anonymous Blocklisting across Service Providers and its Application to Group Messaging

### v1

@@ 4.4.3 Non-frameability
The Non-frameability security property states that it is impossible to prevent an honest non-blocklisted user from successfully authenticating.
The security game is shown in Algorithm4, and it unfolds as follows: the Challenger initialises the environment and allows the adversary to execute arbitrary queries. Eventually, the adversary selects one of the registered users, a target realm and a set of trusted realms. The Challenger then attempts to authenticate the user. The adversary wins if the user has not been banned but the authentication fails.
As shown in Algorithm 1, the $O_{Block}$ query is different for the Non-frameability security game, as it also requires the adversary to submit an user’s secret value $x$ [alpaca]. This allows the Challenger to identify which specific user is being targeted in order to later check if the target user has been blocklisted.
@@ 5 Proposed FAB Scheme
We now provide a concrete construction of a FAB scheme. The proposed FAB scheme uses as primitives a pseudo-random function $PRF$, a hash function $H$, a signature scheme $S=\{\mathsf{S.Sign},\mathsf{S.Verify}\}$, a negative accumulator $AC=\{\mathsf{AC.Create},\mathsf{AC.Add},\mathsf{AC.Remove},\mathsf{AC.NonMemProve},\allowbreak\mathsf{AC.VerNonMem}\}$ and a zk-SNARK scheme $ZK=\{\mathsf{ZK.Gen},\mathsf{ZK.Sign},\mathsf{ZK.Verify}\}$.
Algorithm 5 Initialisation Functions of the FAB scheme.
@@
Algorithm 6 Realm functions
@@
Algorithm 7 Authentication and Verification Functions.
@@
Algorithm 8 Identity Provider Functions
@@
The instantiations of algorithms $\mathsf{Setup}$, $\mathsf{Register}$ and $\mathsf{Create}$ are shown in Algorithm 5, and $\mathsf{Block}$ and $\mathsf{Unblock}$ are shown in Algorithm 6.
The $\mathsf{Auth}$ instantiation is shown in Algorithm 7. It involves the creation of $U$’s pseudonym for target realm $ps^{R}_{U}$ as well as for every trusted realm $ps^{i}_{U}$. Then, a proof of non-membership of the corresponding pseudonym is generated for all realms’ blocklist. A number of zk-SNARK proofs are generated: one for the $R^{auth}$ relation and one for every trusted realm for the $R^{block}$ relation.
The $\mathsf{Verify}$ instantiation, also shown in Algorithm7, follows a similar execution flow. All zk-SNARK proofs generated in the $\mathsf{Auth}$ algorithm are verified, as well as the non-membership proof for the pseudonym $ps^{R}$.
@@
In order to successfully authenticate to a realm, an user must prove (1) that they possess a valid credential signed by the Identity Provider and (2) that they are not banned in any of the trusted realms. We formalise the two statements as the relations $R^{auth}$ and $R^{block}$, to be proven through a zk-SNARK. Figure 3 shows the description of both relations.
@@
The relation $R^{auth}$ is shown in Algorithm 4: its witness is the user’s credential and its statement is the issuer’s public key, the target realm’s seed and the user’s pseudonym in that realm. $R^{auth}$ is satisfied when the pseudonym is correctly computed and the credential’s signature is valid.
@@
$R^{auth}=\left\{\begin{array}[]{@{}l@{}}((pk_{IP},s_{G},ps^{G}_{U}),(\sigma_{U},x_{U}))\end{array}:\begin{array}[]{@{}l@{}}\mathsf{S.Verify_{pk_{IP}}(\sigma_{U},x_{U})},\\ ps^{G}_{U}=H(x_{U},s_{G})\end{array}\right\}$ (4) $R^{block}=\left\{\begin{array}[]{@{}l@{}}((ps^{G}_{U},s_{G},s_{G^{\prime}},acc_{G^{\prime}}),(\overline{w^{G^{\prime}}_{x}},x_{U}))\end{array}:\begin{array}[]{@{}l@{}}ps^{G}_{U}=H(x_{U},s_{G})\\ ps^{G^{\prime}}_{U}=H(x_{U},s_{G^{\prime}})\\ \mathsf{AC.VerNonMem}(acc_{G^{\prime}},ps^{G^{\prime}}_{U},\overline{w^{G^{\prime}}_{x}}),\\ \end{array}\right\}$ (5)
Figure 3: Description of the relations $R^{ban}$ and $R^{block}$.
@@
Algorithm 5 shows the $R^{block}$ relation. The relation mainly checks that the user is not blocked by verifying a non-membership proof of the pseudonym that the user would have in the trusted realm. However, it also checks again the validity of the user’s pseudonym in the target realm: this is necessary to ensure that the same credential is used in both relations.
@@ 5.1 Security
    $G_{1}$ is identical to the Blocklistability game described in Algorithm 3, but $\operatorname{PRF}$ is replaced by a random oracle. As $PRF$ is secure, $G_{1}$ is indistinguishable from the original game.
@@
    $G_{2}$ is identical to $G_{1}$ but the Challenger executes an extractor $E$ for the Auth relation, obtaining $(\sigma,x^{auth})\leftarrow\mathcal{E}(\pi^{auth})$. Since the zk-SNARK scheme is knowledge-sound, the probability that $ps\neq\operatorname{PRF}_{x^{auth}}(s_{R})$ or $\mathsf{S.Verify_{pk_{IP}}}(\sigma,H(x))=Reject$ is negligible.
  • 3.
    $G_{3}$ is identical to $G_{2}$ but the Challenger executes an extractor $E$ for the Ban relation, obtaining $(\overline{w^{i}},x^{block}_{i})\leftarrow\mathcal{E}(\pi^{block}_{i})$ for every $i\in I$. Since the zk-SNARK scheme is knowledge-sound, the probability that $ps\neq\operatorname{PRF}_{x^{block}_{i}}(s_{R})$ or $\mathsf{AC.VerNonMem}(acc_{i},\operatorname{PRF}_{x^{block}_{i}}(s_{i}),\overline{w^{i}})=Reject$ for any $i\in I$ is negligible.
As per $G_{1}$, $\operatorname{PRF}$ has been replaced by a random oracle and thus the chance of finding a collision is negligible. Since $\operatorname{PRF}_{x^{auth}}(s_{R})=PRF_{x^{block}_{i}}(s_{R})=ps$, then with overwhelming probability $x^{auth}=x^{block}_{i}=x$.
Let $B$ be an adversary playing an EUF-CMA game for the signature scheme $S$, that acts as a wrapper for $A$. $B$ will simulate the Challenger for $A$’s game and execute all queries involved in it. During $Q_{Register}$ queries, $B$ requests its challenger for a signature of $H(x)$. Eventually, $A$ outputs $(r,I,ps,\overline{w^{R}},\pi)$ and $B$ obtains $(\sigma,x^{auth})$ from $E$. $B$ then submits $(\sigma,H(x))$ to its challenger. Clearly, if $A$ was able to forge $\sigma$ then $B$ would win its game. Thus, the chance that $A$ successfully uses a $\sigma$ not obtained through the $O_{Register}$ query is negligible. Let credcred be the credential generated through a $Q_{Register}$ query that contains $(\sigma,x)$.
Let $B$ be an adversary playing simultaneous Soundness games for the accumulator $AC$, that acts as a wrapper for $A$. $B$ will simulate the Challenger for $A$’s game and execute all queries involved in it. Whenever $A$ calls $Q_{Create}$, $B$ initiates a game with a new challenger and forwards the received accacc. During $O_{Block}$ and $O_{Unblock}$ queries, $B$ will request the corresponding challenger to add and delete the received $ps$. Eventually, $A$ outputs $(r,I,ps,\overline{w^{R}},\pi)$ and $B$ obtains $\{(\overline{w^{i}},x)\}_{i\in I}$ from $E$. $B$ calculates $\{ps_{i}\leftarrow PRF_{x}(s_{i})\}_{i\in I}$ and submits $(ps,\overline{w^{R}})\cup\{(ps_{i},\overline{w^{i}})\}$ to each corresponding challenger.
Let $j\in I\cup\{r\}$ such that $(j,ps_{j})\in B$ - as specified by the all_blockedall\_blocked predicate, at least one such $j$ exists. That means the $j$-th challenger has been queried with $ps_{j}$. If $A$ is able to produce a $(ps_{j},\overline{w^{j}})$ such that $\mathsf{AC.Verify}(acc_{j},ps_{j},\overline{w^{j}})=Accept$, then $B$ would win its game against the $j$-th challenger. Thus, the probability of $A$ successfully forging a non-membership proof for a blocked pseudonym is negligible. ∎
@@
We start by applying a series of transforms to the security game:
  • 1.
    $G_{1}$ is identical to the Unlinkability game described in Algorithm 3, but $\operatorname{PRF}$ is replaced by a random oracle. As $PRF$ is secure, $G_{1}$ is indistinguishable from the original game.
  • 2.
    $G_{2}$ is identical to $G_{1}$ but every call to $\mathsf{ZK.Prove}(pk,x,w)$ is replaced by $\mathsf{ZK.Sim}(crs,td,x)$. As stated by the Zero-Knowledge property of the zk-SNARK scheme, $G_{2}$ is indistinguishable from $G_{1}$.
Recall that the output of $Q_{Chal}$ is $(ps^{R}_{b},\overline{w^{R}_{b}},(\pi^{auth}_{b},\{\pi^{block}_{i}\}_{i\in I}))$. As per $G_{2}$, none of these values reveals anything about $cred_{b}$: $ps^{R}_{b}$ is indistinguishable from random and $\pi^{auth}_{b}$ and $\{\pi^{block}_{i}\}_{i\in I})$ are simulated proofs. The latter point implies that none of the values of $cred_{b}=(x_{b},\sigma_{b})$ were used in the generation of the proofs. The same logic applies to any $(ps^{R}_{u},\overline{w^{R}_{u}},\pi_{u})$ output by $Q_{Auth}$, as its zk-SNARK proofs are also simulated.
We now show that $A$ cannot link two pseudonyms in different realms to the same $x_{b}$. Let $B$ be an adversary playing a Key Indistinguishability game for the pseudo-random function $\operatorname{PRF}$, that acts as a wrapper for $A$. $B$ will simulate the Challenger for $A$’s game and execute all queries involved in it. Whenever $B$ would execute $PRF$, it instead queries its challenger. $B$ does not generate $cred_{0},cred_{1}$ as they are not needed for the simulated zk-SNARK proofs. Eventually, $A$ outputs $b^{\prime}$ and $B$ forwards it to its challenger. Clearly, if $A$ succeeds, then $B$ would also win the Key Indistinguishability game. As $\operatorname{PRF}$ is assumed to be secure, $A$ cannot obtain any information about the $x_{b}$ used and thus wins with negligible probability. ∎
@@
We start the proof by applying the transforms up to $G_{1}$ defined in the Unlinkability proof to the Non-frameability security game.
The values $(ps^{R}_{u},\overline{w^{R}_{u}},\pi_{u}$) are honestly generated by the Challenger. Thus, the Correctness property of the accumulator ensures that if $ps^{R}_{u}$ has not been added to $acc_{R}$, then $\mathsf{AC.VerNonMem}(acc_{R},ps^{R}_{u},\overline{w^{R}_{u}})$ will accept. Likewise, the Completeness property of the zk-SNARK scheme ensures that if neither $ps^{i}_{u}=\operatorname{PRF}_{cred_{u}.x}(s_{i})$ for $i\in I$ has been added to $acc_{i}$, then $\mathsf{ZK.Verify}(vk^{block},\pi^{block}_{i},(ps^{R}_{u},s_{R},s_{i},acc_{i}))$ will accept. The same applies to the verification of the Auth circuit.
Thus, $A$ can only win by finding and blocking a $ps^{i}_{y}=\operatorname{PRF}_{y}(s_{i})$ s.t. $ps^{i}_{y}=ps^{i}_{u}=PRF_{cred_{u}.x}(s_{i})$, for any $i\in I\cup{r}$. Clearly, one of the checks in $\mathsf{Verify}$ would reject: either $\mathsf{AC.VerNonMem}(acc_{R},ps^{R}_{u},\overline{w^{R}_{u}})$ - if $i=r$ - or $\mathsf{ZK.Verify}(vk^{block},\pi^{block}_{i},(ps^{R}_{u},s_{R},s_{i},acc_{i}))$ - if $i\in I$. However, as per $G_{1}$, $PRF$ has been replaced by a random oracle and thus the chance of finding a collision is negligible. ∎

### v2

@@ 4.4.3 Non-frameability
The Non-frameability property states that it is impossible to prevent a non-blocklisted user from authenticating.
The security game is shown in Figure 6, and it unfolds as follows: the Challenger initialises the environment and allows the adversary to execute arbitrary queries. Eventually, the adversary selects one of the registered users, a target realm and a set of trusted realms. The Challenger then attempts to authenticate the user. The adversary wins if the user has not been blocked but the authentication fails.
As shown in Figure 3, the $O_{Block}$ query is different for the Non-frameability security game, as it also requires the adversary to submit an user’s secret value $x$ [20]. This allows the Challenger to identify which user is being targeted in order to later check if the target user has been blocklisted.
@@ 5 Proposed FAB Scheme
We now provide a concrete construction of a FAB scheme. The proposed FAB scheme uses as primitives a pseudo-random function $PRF$, a hash function $H$, a signature scheme $S=\{\mathsf{S.KeyGen},\mathsf{S.Sign},\mathsf{S.Verify}\}$, a negative accumulator $AC=\{\mathsf{AC.Create},\mathsf{AC.Add},\mathsf{AC.Remove},\mathsf{AC.NonMemProve},\allowbreak\mathsf{AC.VerNonMem}\}$ and a zk-SNARK scheme $ZK=\{\mathsf{ZK.Gen},\mathsf{ZK.Sign},\mathsf{ZK.Verify}\}$. We refer to a user as $U$, a target realm as $R$ and a trusted realm as $R^{\prime}$.
@@
Figure 7: Initialisation Functions of the FAB scheme.
@@
Figure 8: Realm functions.
@@
Figure 9: Authentication and Verification Functions.
@@
Figure 10: Identity Provider Functions.
State. New realms and users are respectively initialised through the algorithms $\mathsf{Register}$ and $\mathsf{Create}$, shown in Figure 7. The realms’s state $st_{R}$ is composed of two elements: a seed $s_{R}$ employed by users to create their pseudonym, and a negative accumulator $acc_{R}$ representing the realm’s blocklist. The users’ state consists on a credential $cred_{U}$ that contains the user’s identity $x$ and $IP$’s signature over $H(x)$.
The instantiation of algorithms $\mathsf{Block}$ and $\mathsf{Unblock}$ are shown in Figure 8. They are simply tasked with adding or removing elements from the realm’s accumulator.
@@
zk-SNARK Relations. In order to successfully authenticate to a realm, an user must prove (1) that they possess a valid credential signed by the Identity Provider and (2) that they are not blocked in any of the trusted realms. We respectively formalise the two statements as the relations $R^{auth}$ and $R^{block}$, to be proven through a zk-SNARK. Figure 11 shows the description of both relations.
@@
The relation $R^{auth}$ is shown in Equation 4: its statement is the issuer’s public key $pk_{IP}$, the target realm’s seed $s_{R}$ and the user’s pseudonym in the target realm $ps^{R}_{U}$, and its witness is the user’s credential $(x_{U},\sigma_{U})$. $R^{auth}$ verifies that the user’s pseudonym is correctly computed from their identity and the realm’s seed, and that the user’s credential is adequately signed by the Identity Provider. We refer to the zk-SNARK proof that proves the $R^{auth}$ relation as $\pi^{auth}$.
@@
$R^{auth}=\left\{\begin{array}[]{l}\begin{array}[]{@{}l@{}}((pk_{IP},s_{R},ps^{R}_{U}),(x_{U},\sigma_{U}))\end{array}\\ \hline\cr\begin{array}[]{@{}l@{}}\mathsf{S.Verify_{pk_{IP}}(\sigma_{U},x_{U})},\\ ps^{R}_{U}=PRF_{x_{U}}(s_{R})\end{array}\end{array}\right\}$ (4) $R^{block}=\left\{\begin{array}[]{l}\begin{array}[]{@{}l@{}}((ps^{R}_{U},s_{R},s_{R^{\prime}},acc_{R^{\prime}}),(\overline{w^{R^{\prime}}_{x}},x_{U}))\end{array}\\ \hline\cr\begin{array}[]{@{}l@{}}ps^{R}_{U}=PRF_{x_{U}}(s_{R})\\ ps^{R^{\prime}}_{U}=PRF_{x_{U}}(s_{R^{\prime}})\\ \mathsf{AC.VerNonMem}(acc_{R^{\prime}},ps^{R^{\prime}}_{U},\overline{w^{R^{\prime}}_{x}}),\\ \end{array}\end{array}\right\}$ (5)
Figure 11: Description of the relations $R^{auth}$ and $R^{block}$.
@@
Equation 5 shows the $R^{block}$ relation. Its statement is the user’s pseudonym in the target realm $ps^{R}_{U}$, the seeds of both the target and trusted realm $(s_{R},s_{R^{\prime}})$, and the accumulator of the target realm $acc_{R^{\prime}}$. Its witness is the user’s identity $x_{U}$ and a proof of non-membership to the target realm’s accumulator $\overline{w^{R^{\prime}}_{x}}$. The relation mainly checks that the user is not blocked in the trusted realm: to that end, the pseudonym the user would have in the trusted realm $ps^{R^{\prime}}_{U}$ is computed. Then, the proof of non-membership for that pseudonym is verified. We note that the $R^{block}$ relation also verifies that the user’s pseudonym fo the target realm is computed correctly. This may seem redundant, as $R^{auth}$ also verified the same condition. However, it is necessary to ensure that the same identity $x_{U}$ was used for both relations. We refer to the zk-SNARK proof that proves the $R^{block}$ relation as $\pi^{block}$.
The $\mathsf{Setup}$ algorithm, shown in Figure 7. $\mathsf{Setup}$ is executed by a trusted party such as the Identity Provider, and it initialises the zk-SNARK public parameters for both relations.
The $\mathsf{Auth}$ instantiation is shown in Figure 9. It involves the creation of $U$’s pseudonym for target realm $ps^{R}_{U}$ as well as for every trusted realm $ps^{i}_{U}$. A zk-SNARK proof for the $R^{auth}$ is created using the target realm’s seed. Then, a zk-SNARK proof of the $R^{block}$ relation is generated for all trusted realms’ blocklists, using the corresponding user pseudonym. Thus, a total of $i+1$ zk-SNARK proofs are generated, where $i$ is the number of trusted realms.
The $\mathsf{Verify}$ instantiation, also shown in Figure9, follows a similar execution flow. All zk-SNARK proofs generated in the $\mathsf{Auth}$ algorithm are verified, as well as the non-membership proof for the pseudonym $ps^{R}$.
@@ 5.1 Security
    $G_{1}$ is identical to the Blocklistability game described in Figure 5, but the Challenger executes an extractor $E$ for the Auth relation, obtaining $(\sigma,x^{auth})\leftarrow\mathcal{E}(\pi^{auth})$. Since the zk-SNARK scheme is knowledge-sound, the probability that $ps\neq\operatorname{PRF}_{x^{auth}}(s_{R})$ or $\mathsf{S.Verify_{pk_{IP}}}(\sigma,H(x))=Reject$ is negligible.
@@
    $G_{2}$ is identical to $G_{1}$ but the Challenger executes an extractor $E$ for the $R^{block}$ relation, obtaining $(\overline{w^{i}},x^{block}_{i})\leftarrow\mathcal{E}(\pi^{block}_{i})$ for every $i\in I$. Since the zk-SNARK scheme is knowledge-sound, the probability that $ps\neq\operatorname{PRF}_{x^{block}_{i}}(s_{R})$ or $\mathsf{AC.VerNonMem}(acc_{i},\operatorname{PRF}_{x^{block}_{i}}(s_{i}),\overline{w^{i}})=Reject$ for any $i\in I$ is negligible.
We have established that $\operatorname{PRF}_{x^{auth}}(s_{R})=PRF_{x^{block}_{i}}(s_{R})=ps$. Since $PRF$ is collision-resistant, the probability of $x^{auth}\neq x^{block}_{i}$ is negligible. Thus, we will refer to them simply as $x$.
Let $B$ be an adversary playing an EUF-CMA game for the signature scheme $S$, that acts as a wrapper for $A$. During $Q_{Register}$ queries, $B$ requests its challenger for a signature of $H(x)$. Eventually, $A$ outputs $(r,I,ps,\overline{w^{R}},\pi)$ and $B$ obtains $(\sigma,x^{auth})$ from $E$. $B$ then submits $(\sigma,H(x))$ to its challenger. Clearly, if $A$ was able to forge $\sigma$ then $B$ would win its game. Thus, the chance that $A$ successfully uses a $\sigma$ not obtained through the $O_{Register}$ query is negligible. Let credcred be the credential generated through a $Q_{Register}$ query that contains $(\sigma,x)$.
Let $B$ be an adversary playing simultaneous Soundness games for the accumulator $AC$, that acts as a wrapper for $A$. Whenever $A$ calls $Q_{Create}$, $B$ initiates a game with a new challenger and forwards the received accacc. During $O_{Block}$ and $O_{Unblock}$ queries, $B$ will request the corresponding challenger to add and delete the received $ps$. Eventually, $A$ outputs $(r,I,ps,\overline{w^{R}},\pi)$ and $B$ obtains $\{(\overline{w^{i}},x)\}_{i\in I}$ from $E$. $B$ calculates $\{ps_{i}\leftarrow PRF_{x}(s_{i})\}_{i\in I}$ and submits $(ps,\overline{w^{R}})\cup\{(ps_{i},\overline{w^{i}})\}$ to each corresponding challenger.
Let $j\in I\cup\{r\}$ such that $(j,ps_{j})\in B$ — as specified by the all_blockedall\_blocked predicate, at least one such $j$ exists. That means the $j$-th challenger has been queried with $ps_{j}$. If $A$ is able to produce a $(ps_{j},\overline{w^{j}})$ such that $\mathsf{AC.Verify}(acc_{j},ps_{j},\overline{w^{j}})=Accept$, then $B$ would win its game against the $j$-th challenger. Thus, the probability of $A$ successfully forging a non-membership proof for a blocked pseudonym is negligible. ∎
@@
We start by applying the following transform to the security game: $G_{1}$ is identical to the Unlinkability game described in Figure 5, but every call to $\mathsf{ZK.Prove}(pk,x,w)$ is replaced by $\mathsf{ZK.Sim}(crs,td,x)$. As stated by the Zero-Knowledge property of the zk-SNARK scheme, $G_{2}$ is indistinguishable from the original game.
Recall that the output of $Q_{Chal}$ is $(ps^{R}_{b},\overline{w^{R}_{b}},(\pi^{auth}_{b},\{\pi^{block}_{i}\}_{i\in I}))$. As per $G_{1}$, none of these values reveals anything about $cred_{b}$: $ps^{R}_{b}$ is indistinguishable from random and $\pi^{auth}_{b}$ and $\{\pi^{block}_{i}\}_{i\in I})$ are simulated proofs. The latter point implies that none of the values of $cred_{b}=(x_{b},\sigma_{b})$ were used in the generation of the proofs. The same logic applies to any $(ps^{R}_{u},\overline{w^{R}_{u}},\pi_{u})$ output by $Q_{Auth}$, as its zk-SNARK proofs are also simulated.
We now show that $A$ cannot link two pseudonyms in different realms to the same $x_{b}$. Let $B$ be an adversary playing a Key Indistinguishability game for the pseudo-random function $\operatorname{PRF}$, that acts as a wrapper for $A$. Whenever $B$ would execute $PRF$, it instead queries its challenger. $B$ does not generate $cred_{0},cred_{1}$ as they are not needed for the simulated zk-SNARK proofs. Eventually, $A$ outputs $b^{\prime}$ and $B$ forwards it to its challenger. Clearly, if $A$ succeeds, then $B$ would also win the Key Indistinguishability game. As $\operatorname{PRF}$ is assumed to be secure, $A$ cannot obtain any information about the $x_{b}$ used and thus wins with negligible probability. ∎
@@
The values $(ps^{R}_{u},\overline{w^{R}_{u}},\pi_{u}$) are honestly generated by the Challenger. Thus, the Correctness property of the accumulator ensures that if $ps^{R}_{u}$ has not been added to $acc_{R}$, then $\mathsf{AC.VerNonMem}(acc_{R},ps^{R}_{u},\overline{w^{R}_{u}})$ will accept. Likewise, the Completeness property of the zk-SNARK scheme ensures that if neither $ps^{i}_{u}=\operatorname{PRF}_{cred_{u}.x}(s_{i})$ for $i\in I$ has been added to $acc_{i}$, then $\mathsf{ZK.Verify}(vk^{block},\pi^{block}_{i},(ps^{R}_{u},s_{R},s_{i},acc_{i}))$ will accept. The same applies to the verification of $R^{auth}$.
Thus, $A$ can only win by finding and blocking a $ps^{i}_{y}=\operatorname{PRF}_{y}(s_{i})$ s.t. $ps^{i}_{y}=ps^{i}_{u}=PRF_{cred_{u}.x}(s_{i})$, for any $i\in I\cup{r}$. Clearly, one of the checks in $\mathsf{Verify}$ would reject: either $\mathsf{AC.VerNonMem}(acc_{R},ps^{R}_{u},\overline{w^{R}_{u}})$ - if $i=r$ - or $\mathsf{ZK.Verify}(vk^{block},\pi^{block}_{i},(ps^{R}_{u},s_{R},s_{i},acc_{i}))$ - if $i\in I$. However, $PRF$ is collision-resistant and thus the chance of finding a collision is negligible. ∎